OpenStudy (anonymous):

EASY MEDAL JUST NEED FAST ANSWERS Write the equation y - 2 = 4(x + 5) in standard form. (Points : 1) 4x + y = 7 4x - y = 3 4x - y = -22

OpenStudy (anonymous):

HELP

OpenStudy (judygreeneyes):

y - 2 = 4(x+5) y = 4X + 20+2 y = 4X + 22 4X - Y = -22

OpenStudy (judygreeneyes):

It was a bit hard telling that the equations you gave after your question were possible answers, but looks like I matched one!